OpenStudy (anonymous):

Solve for x 4x^-3+2=10

OpenStudy (anonymous):

isolate x first

OpenStudy (anonymous):

Then I got 4x^-3=8

OpenStudy (anonymous):

now what does 4x^-3 the same as?

OpenStudy (anonymous):

divide by both sides by 4 and you get x^-3 and (8/4) and then I stuck

OpenStudy (anonymous):

what does x^-3 the same as?

OpenStudy (anonymous):

.001

OpenStudy (anonymous):

no do you know what x^-1 is?

OpenStudy (anonymous):

.1

OpenStudy (anonymous):

no you do not know what x is. what is it in terms of x?

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

You have to solve for x where do you get 1?

OpenStudy (anonymous):

I don't know I am really confused

OpenStudy (anonymous):

x^-1 = 1/x and x^-2 = 1/x^2 so what is x^-3?

OpenStudy (anonymous):

it equals 1

OpenStudy (anonymous):

1/x

OpenStudy (anonymous):

1/x^3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now solve for x by 1/x^3 = 4

OpenStudy (anonymous):

= 2 i mean

OpenStudy (anonymous):

not 4

OpenStudy (anonymous):

ok so then x=8

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

1/x^3 = 2 solve for x

OpenStudy (anonymous):

2.001

OpenStudy (anonymous):

do you know how to cross multiply?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do it to 1/x^3 = 2

OpenStudy (anonymous):

so then I get 2=x^3 then what

OpenStudy (anonymous):

that is not right

OpenStudy (anonymous):

1=2x^3

OpenStudy (anonymous):



OpenStudy (anonymous):

yeah then what

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

isolate x first

OpenStudy (anonymous):

does x=5

OpenStudy (anonymous):

no

OpenStudy (anonymous):